OpenStudy (anonymous):

Find the value of two numbers if their sum is 38 and their difference is 8. Please show work so I can see how its done. Thanks!

OpenStudy (anonymous):

Two numbers a and b, let's say a is larger. a+b=38 a-b=8 Add two equations 2a=46 a=23 Plug this in. b=15 23 and 15

OpenStudy (anonymous):

a+b=38 a-b=8

OpenStudy (anonymous):

wow, you guys are like lightning. Nice job! Thank you very much.

OpenStudy (anonymous):

i would do neither, but instead guess and check

OpenStudy (anonymous):

Hey satellite! I appreciate your advice, do you have any suggestions for me to do this right? Could you show me so I learn correctly?